the diameter of an automobil tyre is normally distributed with a mean of 575mm and a standard deviation of 5mm
what is the probability that a randomly selected tyre has a diameter
a. between 575 and 579mm
b.above 570mm
c.below 580mm
2.80% of the tyre will have diameter above what value

Answer by ewatrrr(24785)


Hi,
using TI function normalcdf(smaller, larger, µ, σ):
mean of 575mm and a standard deviation of 5mm
P(575 ≤ x ≤ 579)= normalcdf(575, 579, 575, 5) 0r using z-value NORMSDIST(.8) - .5
P(x > 570) = normalcdf(570, 9999, 575, 5) 0r using z-value 1 - NORMSDIST(-1.0)
P(x < 580) = normalcdf(-9999, 580, 575, 5) 0r using z-value NORMSDIST(1.0)
2.80% of the tyre will have diameter above what value
z = |NORMSINV(.20/2)| = 1.2816
1.2816 = (X - 575)/5
5*1.2816 + 575 = 581.4mm